# Tensor Product Attention Is All You Need

### v4

@@ Abstract
Scaling language models to handle longer input sequences typically necessitates large key-value (KV) caches, resulting in substantial memory overhead during inference. In this paper, we propose Tensor Product Attention (TPA), a novel attention mechanism that uses tensor decompositions to represent queries, keys, and values compactly, substantially shrinking the KV cache size at inference time. By factorizing these representations into contextual low-rank components and seamlessly integrating with Rotary Position Embedding (RoPE), TPA achieves improved model quality alongside memory efficiency. Based on TPA, we introduce the Tensor ProducT ATTenTion Transformer (T6), a new model architecture for sequence modeling. Through extensive empirical evaluation on language modeling tasks, we demonstrate that T6 surpasses or matches the performance of standard Transformer baselines, including Multi-Head Attention (MHA), Multi-Query Attention (MQA), Grouped-Query Attention (GQA), and Multi-Head Latent Attention (MLA) across various metrics, including perplexity and a range of established evaluation benchmarks. Notably, TPA’s memory efficiency and computational efficiency at the decoding stage enable processing longer sequences under fixed resource constraints, addressing a critical scalability challenge in modern language models. The code is available at https://github.com/tensorgi/T6.
@@ 1 Introduction
A key advantage of TPA is its native compatibility with rotary positional embeddings (RoPE) (su2024roformer, ), enabling a straightforward drop-in replacement for multi-head attention (MHA) layers in modern LLM architectures such as LLaMA (touvron2023llama, ) and Gemma (team2024gemma, ).
@@ 2 Background
In this section, we briefly review Scaled Dot-Product Attention, Multi-Head Attention (vaswani2017attention, ) and introduce key notations. Other attention mechanisms like Multi-Query Attention (MQA) (shazeer2019fast, ), Grouped Query Attention (GQA) (ainslie2023gqa, ), Multi-head Latent Attention (MLA) (liu2024deepseek, ; liu2024deepseekv3, ), and Rotary Position Embedding (RoPE) (su2024roformer, ) are further discussed in Appendix H.
@@ 2.2 Multi-Head Attention (MHA)
MHA enables the model to capture a rich set of dependencies by allowing each head to focus on different aspects of the input sequence. We also discuss how MHA relates to TPA in Appendix A. Further discussion on Transformers and attention mechanisms, KV cache optimization techniques, and low-rank factorization methods is available in Appendix G.
@@ 3.1 Tensor Factorization of Queries, Keys, and Values
[…]
@@
where $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h}$, $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{h}}$. Hence, for queries, each tensor product $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\colon\mathbb{R}^{h}\times\mathbb{R}^{d_{h}}\to\mathbb{R}^{h\times d_{h}}$ (an outer product) contributes to the query slice $\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}$. Analogous definitions apply to the key slice $\mathbf{K}_{t}$ and value slice $\mathbf{V}_{t}$.
@@
Parameter Initialization. We use Xavier initialization (glorot2010understanding, ) for the factor weight matrices; details are in Appendix I.
@@ 3.2 RoPE Compatibility and Acceleration
where $\operatorname{RoPE}_{t-s}$ denotes applying RoPE with relative position $t-s$. Focusing on individual heads $i$, if $\mathbf{q}_{t,i}$ and $\mathbf{k}_{s,i}$ are the $i$-th head vectors from $\mathbf{Q}_{t}$ and $\mathbf{K}_{s}$ respectively (as column vectors of dimension $d_{h}$), and $\widetilde{\mathbf{q}}_{t,i}=\operatorname{RoPE}(\mathbf{q}_{t,i},t)$ and $\widetilde{\mathbf{k}}_{s,i}=\operatorname{RoPE}(\mathbf{k}_{s,i},s)$, then the dot product for attention scores satisfies:
@@
Theorem 3.1 indicates that TPA does not break RoPE’s relative translational property. We prove Theorem 3.1 in Appendix F.1. In essence, $\operatorname{RoPE}_{t}$ applies a linear transformation (rotation matrix $\mathbf{R}_{t}$) to the $d_{h}$-dimensional space. Since $\mathbf{Q}_{t}=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\mathbf{B}_{Q}(\mathbf{x}_{t})$, applying RoPE yields $\mathbf{Q}_{t}\mathbf{R}_{t}=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}(\mathbf{B}_{Q}(\mathbf{x}_{t})\mathbf{R}_{t})$. Thus, RoPE effectively transforms $\mathbf{B}_{Q}(\mathbf{x}_{t})$ to $\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t})=\mathbf{B}_{Q}(\mathbf{x}_{t})\mathbf{R}_{t}$, where $\mathbf{R}_{t}$ acts on each row of $\mathbf{B}_{Q}(\mathbf{x}_{t})$. The $\mathbf{A}_{Q}(\mathbf{x}_{t})$ factor remains unchanged, preserving the TPA structure.
@@ 3.3 KV Caching and Memory Reduction
Table 1: Comparison of different attention mechanisms. Here, $R_{Q}$, $R_{K}$, and $R_{V}$ denote the ranks for queries, keys, and values in TPA, respectively. Variants of TPA, such as TPA (KVonly), TPA (Non-contextual A), and TPA (Non-contextual B), are detailed in Appendix I. For MLA, $d_{h}^{R}$ and $d_{h}$ are the dimensions for RoPE and non-RoPE parts; $d_{c}^{\prime}$ and $d_{c}$ are the dimensions of compressed vectors for query and key-value, respectively.
@@
MLA $d_{c}+d_{h}^{R}$ $d_{c}^{\prime}(d_{\text{model}}+hd_{h}+hd_{h}^{R})$ $+d_{\text{model}}d_{h}^{R}+d_{c}(d_{\text{model}}+2hd_{h})$ $h$ $h$
@@
3.4 Expressing MHA, MQA, and GQA as Non-contextual TPA
We demonstrate in Appendix A that several existing attention mechanisms, including MHA, MQA, and GQA, can be elegantly expressed as non-contextual variants of TPA by imposing specific constraints on the TPA factors.
@@ 3.5 Model Architectures
We propose a new architecture called Tensor ProducT ATTenTion Transformer (T6), which uses our Tensor Product Attention (TPA) in place of standard MHA (multi-head attention) or GQA (grouped-query attention). Building upon the query, key, and value tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ defined in Section 3.1, T6 utilize the overall architecture of LLaMA (touvron2023llama, ) while changing the self-attention block to our TPA-based version. The feed-forward network (FFN) adopts a SwiGLU layer, as in (shazeer2020glu, ; touvron2023llama, ).
The specific mechanisms for TPA-based query, key, and value factorization, as well as the attention computation within T6, directly follow the comprehensive descriptions provided in Section 3 (particularly Section 3.1 for factorization, and Equations 3.2 and 3.3 for attention computation). A brief recap of these steps within the T6 model context is provided in Appendix B.
Rotary Positional Embedding (RoPE). As discussed in Section 3.2, RoPE (su2024roformer, ) is applied to the $\mathbf{Q}$ and $\mathbf{K}$. Within TPA, we pre-rotate the factor $\mathbf{b}^{Q}_{t}(\mathbf{x}_{t})$ and $\mathbf{b}^{K}_{s}(\mathbf{x}_{s})$ directly, so that each $\mathbf{K}_{s}$ is already rotated prior to caching, see (3.5) and Theorem 3.1.
@@ 4 FlashTPA Decoding Algorithm
For efficient autoregressive inference with Tensor Product Attention (TPA), we introduce FlashTPA Decoding. This algorithm is optimized for generating one token at a time by leveraging the factorized representation of queries, keys, and values. The core idea, illustrated in Figure 2, is to perform attention computations using a sequence of Einstein summations (“einsum”) that operate directly on these factorized components. This avoids materializing the full query, key, and value tensors, which is particularly beneficial as the Key-Value (KV) cache grows with sequence length.
The detailed definitions of the input factorized components and the step-by-step pseudocode for FlashTPA Decoding are provided in Algorithm 1 (see Appendix D for details). An optimized Triton kernel implementation: Algorithm 3 is also outlined in Appendix D.
This sequence of factorized operations allows FlashTPA Decoding to compute the attention output efficiently. It minimizes memory allocations for large intermediate tensors and reduces the overall computational load compared to materializing full query, key, and value tensors, especially for long sequences. Consequently, TPA is not only memory-efficient due to its smaller KV cache footprint but can also be computationally efficient during inference. The experimental results for FlashTPA decoding time are presented in Section 5.2.
Algorithm 1 FlashTPA Decoding Algorithm
1:$\mathbf{A}_{Q}$: Query A tensor $(B,1,H,R_{Q})$
2:$\mathbf{B}_{Q}$: Query B tensor $(B,1,R_{Q},D)$
3:$\mathbf{a}^{K}_{\text{cache}}$: Cached Key A tensor $(B,M,H)$
4:$\bm{b}^{K}_{\text{cache}}$: Cached Key B tensor $(B,M,D)$
5:$\mathbf{a}^{V}_{\text{cache}}$: Cached Value A tensor $(B,M,H)$
6:$\bm{b}^{V}_{\text{cache}}$: Cached Value B tensor $(B,M,E)$
7:$s_{\text{total}},s_{Q},s_{K},s_{V}$: Scaling factors (with defaults $s_{\text{total}}\leftarrow 1/\sqrt{D}$, $s_{Q}\leftarrow 1/R_{Q}$)
8:$\mathbf{O}$: Output tensor $(B,1,H,E)$
9$\triangleright$ Step 1: Interaction between $\mathbf{B}_{Q}$ and $\bm{b}^{K}_{\text{cache}}$ components
10:$S^{(1)}\leftarrow\text{einsum}(\text{``bnrd, bmd}\rightarrow\text{bnmr''},\mathbf{B}_{Q},\bm{b}^{K}_{\text{cache}})$triangleright Shape: $(B,1,M,R_{Q})$. $S^{(1)}_{b,n,m,r}=\sum_{d}(\mathbf{B}_{Q})_{b,n,r,d}\cdot(\bm{b}^{K}_{\text{cache}})_{b,m,d}$
11$\triangleright$ Step 2: Incorporate $\mathbf{A}_{Q}$ component
12:$S^{(2)}\leftarrow\text{einsum}(\text{``bnhr, bnmr}\rightarrow\text{bnmh''},\mathbf{A}_{Q},S^{(1)})$triangleright Shape: $(B,1,M,H)$. $S^{(2)}_{b,n,m,h}=\sum_{r}(\mathbf{A}_{Q})_{b,n,h,r}\cdot S^{(1)}_{b,n,m,r}$
13$\triangleright$ Step 3: Incorporate $\mathbf{a}^{K}_{\text{cache}}$ component to get full logits
14:einsum(“bnmh, $\mathcal{L}\leftarrow\text{einsum}(\text{``bnmh, bmh}\rightarrow\text{bhnm''},S^{(2)},\mathbf{a}^{K}_{\text{cache}})$triangleright Shape: $(B,H,1,M)$. $\mathcal{L}_{b,h,n,m}=S^{(2)}_{b,n,m,h}\cdot(\mathbf{a}^{K}_{\text{cache}})_{b,m,h}$
15$\triangleright$ Step 4: Apply scaling and Softmax
16:$\bm{\alpha}\leftarrow\operatorname{Softmax}_{M}(\mathcal{L}\cdot s_{Q}\cdot s_{K}\cdot s_{\text{total}})$triangleright Shape: $(B,H,1,M)$. Softmax over cache dimension $M$
17$\triangleright$ Step 5: Compute weighted sum with $\mathbf{a}^{V}_{\text{cache}}$ component
18:$\mathbf{O}^{(A)}\leftarrow\text{einsum}(\text{``bhnm, bmh}\rightarrow\text{bnmh''},\bm{\alpha},\mathbf{a}^{V}_{\text{cache}})$triangleright Shape: $(B,1,M,H)$. $\mathbf{O}^{(A)}_{b,n,m,h}=\bm{\alpha}_{b,h,n,m}\cdot(\mathbf{a}^{V}_{\text{cache}})_{b,m,h}$
19$\triangleright$ Step 6: Incorporate $\bm{b}^{V}_{\text{cache}}$ component and final scaling
[…]
21:return $\mathbf{O}$

### v5

@@ Abstract
Scaling language models to handle longer input sequences typically necessitates large key-value (KV) caches, resulting in substantial memory overhead during inference. In this paper, we propose Tensor Product Attention (TPA), a novel attention mechanism that uses tensor decompositions to represent queries, keys, and values compactly, substantially shrinking the KV cache size at inference time. By factorizing these representations into contextual low-rank components and seamlessly integrating with Rotary Position Embedding (RoPE), TPA achieves improved model quality alongside memory efficiency. Based on TPA, we introduce the Tensor ProducT ATTenTion Transformer (T6), a new model architecture for sequence modeling. Through extensive empirical evaluation on language modeling tasks, we demonstrate that T6 surpasses or matches the performance of standard Transformer baselines, including Multi-Head Attention (MHA), Multi-Query Attention (MQA), Grouped-Query Attention (GQA), and Multi-Head Latent Attention (MLA) across various metrics, including perplexity and a range of established evaluation benchmarks. Notably, TPA’s memory efficiency and computational efficiency at the decoding stage enable processing longer sequences under fixed resource constraints, addressing a critical scalability challenge in modern language models. Project Page: https://github.com/tensorgi/TPA.
@@ 1 Introduction
A key advantage of TPA is its native compatibility with rotary positional embeddings (RoPE) (su2024roformer, ), enabling a straightforward drop-in replacement for multi-head attention (MHA) layers in modern LLM architectures such as LLaMA (touvron2023llama, ), Qwen (bai2023qwen, ), and Gemma (team2024gemma, ).
@@ 2 Background
In this section, we briefly review Scaled Dot-Product Attention, Multi-Head Attention (vaswani2017attention, ), and introduce key notations. Other attention mechanisms like Multi-Query Attention (MQA) (shazeer2019fast, ), Grouped Query Attention (GQA) (ainslie2023gqa, ), Multi-head Latent Attention (MLA) (liu2024deepseek, ; liu2024deepseekv3, ), and Rotary Position Embedding (RoPE) (su2024roformer, ) are further discussed in the Appendix F.
@@ 2.2 Multi-Head Attention (MHA)
MHA enables the model to capture a rich set of dependencies by allowing each head to focus on different aspects of the input sequence. We also discuss how MHA, MQA, and GQA relate to TPA in the Section 4.
@@ 3.1 Tensor Factorization of Queries, Keys, and Values
$\mathbf{Q}_{t}=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\hskip 21.52771pt\mathbf{K}_{t}=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),$
@@
where $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h}$$\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{h}}$. Hence, for queries, each tensor product $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\colon\mathbb{R}^{h}\times\mathbb{R}^{d_{h}}\to\mathbb{R}^{h\times d_{h}}$ (an outer product) contributes to the query slice $\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}$. Analogous definitions apply to the key slice $\mathbf{K}_{t}$ and value slice $\mathbf{V}_{t}$.
@@
Parameter Initialization. We use Xavier initialization (glorot2010understanding, ) for the factor weight matrices; details are in the Appendix G.
@@ 3.2 RoPE Compatibility and Acceleration
where $\operatorname{RoPE}_{t-s}$ denotes applying RoPE with relative position tst-s. Focusing on individual heads $i$, if $\mathbf{q}_{t,i}$ and $\mathbf{k}_{s,i}$ are the $i$-th head vectors from $\mathbf{Q}_{t}$ and $\mathbf{K}_{s}$ respectively (as column vectors of dimension $d_{h}$), and $\widetilde{\mathbf{q}}_{t,i}=\operatorname{RoPE}(\mathbf{q}_{t,i},t)$ and $\widetilde{\mathbf{k}}_{s,i}=\operatorname{RoPE}(\mathbf{k}_{s,i},s)$, then the dot product for attention scores satisfies:
@@
Theorem 3.1 indicates that TPA does not break RoPE’s relative translational property. We prove it in the Appendix D.1.
@@ 3.3 KV Caching and Memory Reduction
Table 1: Comparison of different attention mechanisms. Here, $R_{Q}$, $R_{K}$, and $R_{V}$ denote the ranks for queries, keys, and values in TPA, respectively. Variants of TPA, such as TPA (KVonly), TPA (Non-contextual A), and TPA (Non-contextual B), are detailed in the Appendix G. For MLA, $d_{h}^{R}$ and $d_{h}$ are the dimensions for RoPE and non-RoPE parts; $d_{c}^{\prime}$ and $d_{c}$ are the dimensions of compressed vectors for query and key-value, respectively.
@@
MLA $d_{c}+d_{h}^{R}$ $d_{c}^{\prime}(d_{\text{model}}+hd_{h}+hd_{h}^{R})$ $d_{\text{model}}d_{h}^{R}+d_{c}(d_{\text{model}}+2hd_{h})$ $h$ $h$
@@
4 Expressing MHA, MQA, GQA as Non-contextual TPA
We demonstrate that standard Multi-Head Attention (MHA), Multi-Query Attention (MQA), and Grouped-Query Attention (GQA) can be expressed as special, non-contextual variants of Tensor Product Attention (TPA). This is achieved by imposing specific constraints on the TPA factors, particularly by making the head-dimension factors ($\mathbf{a}$) independent of the input token ($\mathbf{x}_{t}$).
4.1 MHA as Non-contextual TPA
Standard Multi-Head Attention (MHA) can be precisely formulated as a TPA where the rank is equal to the number of heads ($R_{Q}=R_{K}=R_{V}=h$), and the head-dimension factors are fixed, non-contextual basis vectors. To recover MHA, we set the rank $R_{Q}=h$ and define the factors for each head $i\in[h]$ as follows:
  • Contextual token factor: This is the standard linear projection for the $i$-th head’s query:
    $\mathbf{b}^{Q}_{i}(\mathbf{x}_{t})=(\bm{W}^{Q}_{i})^{\top}\mathbf{x}_{t}\in\mathbb{R}^{d_{h}}$
  • Non-contextual head factor: This factor is a scaled standard basis vector, independent of $\mathbf{x}_{t}$:
    $\mathbf{a}^{Q}_{i}=h\cdot\mathbf{e}_{i}\in\mathbb{R}^{h}$
    where $\mathbf{e}_{i}$ is the $i$-th standard basis vector (a vector of zeros with a one at the $i$-th position).
Substituting these into the TPA equation, the $1/R_{Q}=1/h$ scaling factor cancels with the scaling of the $\mathbf{a}^{Q}_{i}$ factor:
$\mathbf{Q}_{t}$ $=\frac{1}{h}\sum_{i=1}^{h}\left(h\cdot\mathbf{e}_{i}\right)\otimes\left((\bm{W}^{Q}_{i})^{\top}\mathbf{x}_{t}\right)=\sum_{i=1}^{h}\mathbf{e}_{i}\otimes\left((\bm{W}^{Q}_{i})^{\top}\mathbf{x}_{t}\right)$
The resulting tensor product, $\mathbf{e}_{i}\otimes\mathbf{b}^{Q}_{i}(\mathbf{x}_{t})$, produces an $h\times d_{h}$ matrix where only the $i$-th row is non-zero and contains the vector $(\mathbf{b}^{Q}_{i}(\mathbf{x}_{t}))^{\top}$. Summing these matrices for $i=1,\dots,h$ assembles the complete query tensor $\mathbf{Q}_{t}$, where the $i$-th row is precisely the query vector for the $i$-th head in standard MHA. An analogous construction applies to the key ($\mathbf{K}_{t}$) and value ($\mathbf{V}_{t}$) tensors.
Thus, MHA is equivalent to a non-contextual TPA where the head-dimension factors are fixed and orthogonal, effectively assigning a dedicated rank component to each attention head.
4.2 MQA and GQA as Non-contextual TPA
Similarly, Multi-Query Attention (MQA) and Grouped-Query Attention (GQA) can be seen as non-contextual TPAs where the key and value tensors are formed with a rank lower than the number of heads.
  • MQA as Rank-1 TPA (for K and V). In MQA, all $h$ query heads share a single key and value. This corresponds to a TPA with ranks $R_{K}=1$ and $R_{V}=1$. The key tensor $\mathbf{K}_{t}$ is formed using a single, non-contextual head-dimension factor $\mathbf{a}^{K}=\mathbf{1}_{h}$ (a vector of all ones) and a single contextual token-dimension factor $\mathbf{b}^{K}(\mathbf{x}_{t})=(\bm{W}^{K})^{\top}\mathbf{x}_{t}$:
    $\mathbf{K}_{t}=\frac{1}{1}\left(\mathbf{1}_{h}\otimes\mathbf{b}^{K}(\mathbf{x}_{t})\right)$
    This creates an $h\times d_{h}$ matrix where every row is the same shared key vector $(\mathbf{b}^{K}(\mathbf{x}_{t}))^{\top}$. The same logic applies to the value tensor $\mathbf{V}_{t}$. The queries remain full-rank ($R_{Q}=h$) as in MHA.
  • GQA as Rank-G TPA (for K and V). GQA is an intermediate approach where $h$ heads are divided into $G$ groups, with heads in the same group sharing a key and value. This is equivalent to a TPA with ranks $R_{K}=G$ and $R_{V}=G$. The key tensor is formed by summing $G$ components:
    $\mathbf{K}_{t}=\frac{1}{G}\sum_{j=1}^{G}\mathbf{a}^{K}_{j}\otimes\mathbf{b}^{K}_{j}(\mathbf{x}_{t})$
    Here, $\mathbf{b}^{K}_{j}(\mathbf{x}_{t})$ is the shared key vector for group $j$. The non-contextual factor $\mathbf{a}^{K}_{j}$ is a scaled mask vector, defined as $\mathbf{a}^{K}_{j}=G\cdot\text{mask}_{j}$, where the $\text{mask}_{j}$ vector has ones for heads belonging to group $j$ and zeros elsewhere. This scaling cancels the $1/G$ pre-factor:
    $\mathbf{K}_{t}=\frac{1}{G}\sum_{j=1}^{G}(G\cdot\text{mask}_{j})\otimes\mathbf{b}^{K}_{j}(\mathbf{x}_{t})=\sum_{j=1}^{G}\text{mask}_{j}\otimes\mathbf{b}^{K}_{j}(\mathbf{x}_{t})$
    For example, with $h=8$ heads and $G=2$ groups ($2$ KV heads), the factor for the first group of 4 heads would be $\mathbf{a}^{K}_{1}=2\cdot[1,1,1,1,0,0,0,0]^{\top}$. This construction correctly assembles the final key tensor by broadcasting each group’s shared key to its designated heads without any unintended extra scaling.
This perspective highlights that MHA, MQA, and GQA are specific instances of a more general TPA framework, where expressiveness and parameter sharing are controlled by the rank and the nature (contextual vs. non-contextual) of the tensor factors.
@@ 4.3 Model Architectures
We propose a new architecture called Tensor ProducT ATTenTion Transformer (T6), which uses our Tensor Product Attention (TPA) in place of standard MHA (multi-head attention) or GQA (grouped-query attention). Building upon the query, key, and value tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ defined in Section 3.1, T6 utilizes the overall architecture of LLaMA (touvron2023llama, ) while changing the self-attention block to our TPA-based version. The feed-forward network (FFN) adopts a SwiGLU layer, as in (shazeer2020glu, ; touvron2023llama, ).
Rotary Positional Embedding (RoPE). As discussed in Section 3.2, RoPE (su2024roformer, ) is applied to the $\mathbf{Q}$ and $\mathbf{K}$. Within TPA, we pre-rotate the factor $\mathbf{b}^{Q}_{t}(\mathbf{x}_{t})$ and $\mathbf{b}^{K}_{s}(\mathbf{x}_{s})$ directly, so that each $\mathbf{K}_{s}$ is already rotated prior to caching, see Equation (3.5) and Theorem 3.1.
@@ 5 FlashTPA Decoding Algorithm
For efficient autoregressive inference with Tensor Product Attention (TPA), we introduce FlashTPA Decoding. This algorithm is optimized for generating one token at a time by leveraging the factorized representation of queries, keys, and values. The core idea, illustrated in Figure 2, is to perform attention computations using a sequence of Einstein summations (“einsum”) that operate directly on these factorized components. This avoids materializing the full query, key, and value tensors, which is particularly beneficial as the Key-Value (KV) cache grows with sequence length. The detailed definitions of the input factorized components and the step-by-step pseudo-code for FlashTPA Decoding are provided in Algorithm 2. An optimized Triton kernel implementation is outlined inAlgorithm 3 (see Appendix B.1).
@@
This sequence of factorized operations allows FlashTPA Decoding to compute the attention output efficiently. Consequently, TPA is not only memory-efficient due to its smaller KV cache footprint but can also be computationally efficient during inference. The experimental results for FlashTPA decoding time are presented in Section 6.2.